Using the following notations:

θ = independent scalar variable

r(θ) = position vector depending on the variable θ

tˆ = unit tangent vector

nˆ = unit normal vector

show that

nˆ ⊥ tˆ


1
Expert's answer
2021-03-16T11:35:58-0400

Given,

"\theta=" Independent scalar variable

"r(\theta)=" Position vector

"\hat{t}=" unit tangent vector

"\hat{n}=" unit normal vector

"r(\theta)=r\cos\theta \hat{i}+r\sin\theta\hat{j}"

"\hat{n}=\frac{\overrightarrow{r(\theta)}}{|r(\theta)|}=\frac{r\cos\theta \hat{i}+r\sin\theta\hat{j}}{r}"


"=\cos\theta \hat{i}+\sin\theta\hat{j}"


unit tangent vector"(\hat{t})=\frac{r'(\theta)}{|r'|}"


"\hat{t}=\frac{-r\sin\theta\hat{i}+r\cos\theta\hat{j}}{r}=-\sin\theta \hat{i}+r\cos\theta\hat{j}"

Now,

"\hat{n}.\hat{t}=(\cos\theta \hat{i}+\sin\theta\hat{j})(-\sin\theta \hat{i}+r\cos\theta\hat{j})"

"=-\cos\theta . \sin\theta+\sin\theta.\cos\theta"

"=0"

Hence, it is prove that "\hat{n}" is perpendicular to "\hat{t}" .